# When Pre-trained Visual Representations Fall Short: Limitations in Visuo-Motor Robot Learning

### v1

@@ 3.3 Attending to Policy Related Features
We posit that training policies using the global features of PVRs (i.e., CLS token for ViTs or average channel feature for ResNets) leads to overfitting to scene conditions that are irrelevant to the task at hand. The output features of these representations often capture visual characteristics of the scene that may be irrelevant to the policy (e.g., the texture of a tabletop). Processing such extraneous information not only dilutes the policy network’s focus but also leads to overfitting to specific scene conditions.
This observation aligns with recent work on vision model evaluation (Chen et al., 2024), which argues that only specific image regions carry the necessary information for solving a task. Building on this insight, we hypothesise that incorporating local information is particularly effective in the context of robot learning, echoing findings in PVR distillation research (Shang et al., 2024), though this area remains empirically underexplored.
@@
Recognizing the importance of local information is only part of the solution; a data-driven mechanism is also required to filter irrelevant details, such as background patches, and prioritise task-relevant information. To this end, we adopt the attentive probing methodology to implement Attentive Feature Aggregation (AFA). Specifically, we append a cross-attention layer to the frozen PVR, modified to include a trainable query token that interacts with the sequence of local tokens produced by the model. These tokens correspond to the per-patch embeddings for ViTs and the channel embeddings for ResNets, both from the final layer.
@@
Following Eq. 3, the query token $q$ computes dot products with the feature sequence, with length equal to #patches and dimension $d_{k}$, organized as a matrix $F$. These dot products are passed through a softmax function to assign weights to the contributions of each local token to the final embedding. Also, our module consists of multiple heads, so that specific dimension groups that might be irrelevant to the policy can be filtered out. Gradients are allowed to flow through the cross-attention layer, updating the parameters of $q$ as well as the key and value projection matrices, $W_{K}$ and $W_{V}$.
$\text{Attention}(q,F)=\text{softmax}\left(\frac{q\cdot(F\cdot W_{K})^{\top}}{\sqrt{d_{k}}}\right)F\cdot W_{V}$ (3)
@@ 4 Experiments
4.1 Setup
Environment. We conduct our experiments in the widely used MetaWorld simulation environment (Yu et al., 2020), which is built on the MuJoCo (Todorov et al., 2012) physics engine. From this benchmark, we select the ten tasks visualised in Fig. 3 and generate 25 expert demonstrations with 175 rollout steps for each using the provided heuristic policies. The primary criterion for task selection is to maintain a balanced representation of easy, medium, and hard tasks, as identified in prior work on PVR-based visuo-motor control (Mete et al., 2024; Hu et al., 2023), as well as in our empirical results.
PVRs. To validate our hypotheses, we deploy seven Residual Networks (ResNets) (He et al., 2016) and seven Vision Transformers (ViT) (Dosovitskiy et al., 2021), as summarised in Tab. A1 in the Appendix. Our selection includes the most popular PVRs utilised in robot learning applications that have led to SoTA performance. We also focused on ensuring diversity in training strategies, datasets, and the balance between local and global perception. Despite these variations, we maintain a consistent backbone architecture of ResNet-50 or ViT-B/16, with the exception of DINOv2, which employs a smaller patch size of 14. Also, for DINOv2 we discard overlapping patches, to ensure fairness in the comparison of PVRs. The models tested include powerful representations from vision-specific approaches (e.g., DINO), vision-language models (e.g., CLIP), and robot-learning-focused models (e.g., R3M).
Policy training. For all policy training, we train the policy network five times, using five different seeds, keeping the PVR frozen, and report the interquartile mean (IQM) success rate. As is common practice in similar work (Parisi et al., 2022; Nair et al., 2022; Hu et al., 2023), our policy head consists of a small number of MLP layers (four in our case), separated by ReLUs, and outputs the predicted action. We train with mini-batches of 128 samples for 80,000 steps.
Figure 5: Robustness evaluation of policies trained with PVR-extracted features temporally augmented with TE, and with features processed using AFA and temporally augmented with TE. The evaluation is conducted under perturbed lighting conditions and tabletop texture changes. Subplot (a) shows the average performance per task, while subplot (b) shows the average performance per model.
@@ 4.2 Temporal Encoding
We evaluate the performance of a policy network trained for each PVR under three conditions: without any temporal component, using the three most recent past observations and their latent differences (i.e., FLARE (Shang et al., 2021)), and with our proposed Temporal Encoding (TE) method. We select the dimensionality of TE to be $64$ and the scale parameter $100$, after tuning these hyperparameters in a subset of tasks and PVRs (see Appendix C.3).
Fig. 4 illustrates (a) the average performance per task for each of the three approaches and (b) the average performance per model. Statistical analysis (paired t-tests, Wilcoxon tests-results in Appendix C.2) confirms TE’s gains over FLARE and no augmentation are significant. While VC-1 and iBOT achieve slightly higher average scores with FLARE, all other PVRs benefit significantly from temporal augmentation, even when compared to FLARE-augmented results. Similarly, apart from the “Pick and Place” and “Shelf Place” tasks, TE significantly enhances the average task performance. We attribute TE’s superiority over FLARE to the fact that FLARE’s method of stacking sequential latent embeddings, along with their differences, results in embeddings that are nearly identical due to the high similarity of features extracted from consecutive observations. Consequently, the differences between these embeddings are often near-zero vectors, limiting FLARE’s ability to effectively leverage temporal information. Table A2 provides a detailed summary of the performance of each model-task pair.
An intriguing observation is that PVRs pre-trained with a temporal component in their objective function also benefit considerably from TE. For instance, R3M employs time-contrastive learning (Sermanet et al., 2017) to enforce similarity between representations of temporally adjacent frames-experience substantial gains from TE. Notably, VIP achieves an average performance boost of approximately $15$ percentage points, making it one of the most positively impacted models. This finding suggests a potential reconsideration of how temporal perception is integrated into features designed for robot learning. It raises the possibility that existing approaches may not fully exploit the temporal structure necessary for optimal performance.
Figure 6: Ablation study on our approach (i.e., PVR+TE+AFA). Results are reported on in-domain scenes.
@@ 4.3 Attentive Feature Aggregation
We utilise a cross-attention layer and use 12 attention heads for features extracted from ViTs and 32 for features from ResNets. This configuration ensures that we process 64-dimensional feature chunks in both cases, maintaining fairness between the two backbone architectures. Consistent with standard attentive probing methodologies (Chen et al., 2024; Danier et al., 2024), we employ a cosine learning rate scheduler with a warm-up phase. In all experiments, TE is applied to augment features with a temporal component. Specifically, for AFA, the temporal feature is concatenated with the output of the cross-attention module. We validate our full model’s significance via an ablation study, presented in Fig. 6, that evaluates the contributions of TE and AFA in our approach, comparing it against training the policy network solely with PVR-extracted features.
@@
We hypothesise that AFA learns to attend only to scene areas that are important to the task, disregarding features irrelevant to the policy. To properly evaluate this component, we do not limit our evaluation to in-domain scenes but also with environments recreated with scene perturbations, leaving the training dataset distribution unchanged. These perturbations include two modifications, details of which are included in Appendix D.2. First, tabletop texture changes, randomly selected from $30$ distinct textures, some of which feature vibrant patterns that act as strong distractors. These changes affect a significant area of the frame. Second, variations in lighting conditions, including adjustments to the orientation, position, and brightness of the light source. These modifications influence the entire frame, including the robot and the object it manipulates.
We summarize the performance of policies trained directly with PVR features and those trained with features aggregated using AFA in Fig. 5, averaged across tasks and models on perturbed scenes. We also provide in Fig. A4 a concise summary of the performance of each task and model pair on in-domain scenes and on perturbed scenes (per perturbation category).
From these results, several trends emerge. AFA consistently improves the robustness of policies across all tasks, often yielding significant gains, with some tasks showing up to a threefold improvement. The only exception is the “Drawer Open” task, which sees little benefit from AFA. This is likely due to its inherent simplicity, as the task involves manipulating a large object without requiring control of the gripper fingers, which remain open throughout the demonstrations. Consequently, attending to local observations has limited impact.
@@
Additionally, most models exhibit improved out-of-distribution performance with AFA, except for CLIP and ViT-B/16. This is reasonable, as these models are trained with objectives that emphasise global frame perception, unlike other models that incorporate supervision at the patch level. Notably, MIM-trained PVRs benefit the most from AFA, reflecting the alignment between AFA’s design, which is inspired by attentive probing, and the training principles of MIM-based models. These findings highlight AFA’s ability to enhance policy performance, particularly in challenging out-of-distribution scenarios, and underscore its compatibility with models that leverage local feature representations.
@@
The average in-domain performance remains nearly unchanged, with a slight increase from $63.1$% to $66.4$% when using AFA. The minor performance boost observed with AFA in in-domain scenarios, especially when compared to its substantial improvements in perturbed scenes, suggests that AFA does not learn a new latent space for the PVR that is more suited to the task. Instead, it appears to refine the use of the existing latent space by learning to effectively leverage relevant information while discarding elements that are irrelevant to the policy. This distinction underscores AFA’s role as a mechanism for better utilisation of pre-trained features rather than redefining or adapting the underlying feature space.
Figure 7: Comparison of attention heatmaps between the DINO and DINO+AFA (ours). The latter learns to focus on task-relevant regions and ignores scene changes (e.g., distractor objects). Visualisations from more PVRs are included in Fig. A7.
@@ 5 Discussion
We explored how to effectively utilise PVRs for visuo-motor policy learning, identifying the issues of feature temporal entanglement and lack of robustness in scene visual changes. Furthermore, we proposed two approaches to address these limitations leading to a significant performance increase. Below we provide a set of questions with answers that we believe further highlight the contributions of our work.
Q:Does TE help because expert demos are time synced?
A: Even though the expert demos are generated by a heuristic policy, in Fig. A3 we show that they are actually not synchronised and the only common factor is that the task is always completed by T=175 steps.
Q:Does AFA’s robustness stem from the additional trainable parameters?
A: AFA introduces approximately 1.5M additional parameters to a policy model that originally has around 0.5M parameters. To determine whether the observed performance improvement is simply due to the increased model capacity, we train a policy network using raw PVR features but design a deeper policy network to match AFA’s parameter count. The results, shown for two tasks in Fig. A6, suggest otherwise. This indicates that the performance gains arise from AFA’s ability to effectively attend to local information, as is suggested by Fig. 7.
Q:Does the simulation’s lack of visual realism affect the perception of PVRs?
A: Empirical evidence has shown that simulations serve as a reliable proxy for PVR-based visuo-motor policy learning, particularly in environments like MetaWorld (Silwal et al., 2024). To further validate this, we replicate the visualization in Fig. 1 for an expert demonstration in two real-world tasks. The results in Fig. A2 and Fig. A1, reveal that the same feature-related issues persist, underscoring the relevance of our findings across both simulated and real-world settings.
Q:Why is there still a significant performance gap under scene perturbations?
A: We posit that the remaining performance gap largely stems from inherent limitations of behaviour cloning. Its supervised nature makes the policy network vulnerable to compounding errors. Despite improved feature attention, the models tend to overfit to specific visual settings, making it sensitive to domain shifts that progressively push the policy out-of-distribution.
Q:Can we get even better policy performance?
A: Behaviour cloning performance is largely dependent on how well the expert policy $\pi^{\star}$ is sampled. Increasing the number of demonstrations or utilising techniques like DAgger (Ross et al., 2011) could help mitigate this issue.
Limitations. We conducted extensive experiments, evaluating multiple PVRs and training policies across different seeds and representative tasks. However, our work has certain limitations. First, while MetaWorld serves as a reliable and widely used proxy for robot learning experiments, we have not tested our methods on a real robot. Evaluating their effectiveness in real-world scenarios remains an important direction for future research. Second, in line with common practice, we use a simple MLP policy network to ensure our approach targets the feature space and remains agnostic to policy architecture. However, it would be valuable to explore how our methods impact state-of-the-art policy models, such as diffusion-based approaches, which are known to struggle when conditioned on PVRs (Chi et al., 2023). Future work should address these limitations by assessing our approach with alternative policy architectures trained end-to-end with visual encoders.
@@ 6 Conclusion
The use of PVRs for visuo-motor policy learning is still in its early stages, and we believe our work paves the way for further exploration of key challenges. Our insights contribute to the development of PVR models specifically designed for robot learning, ultimately leading to a generalist robotic system powered by large-scale vision foundation models.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning and Robotics. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v2

@@ 3.3 Attending to Policy Relevant Features
We posit that training policies using the global features of PVRs (i.e., CLS token for ViTs or average pooled channel feature for CNNs) can lead to overfitting to scene conditions that are irrelevant to the task at hand. The output features of these representations often capture visual characteristics of the scene that may be irrelevant to the policy (e.g., the texture of a tabletop). Processing such extraneous information not only dilutes the policy network’s focus, but also leads to overfitting to scene specific conditions. This observation aligns with recent work on vision model evaluation [33], which argues that only specific image regions carry the necessary information for solving a task. Building on this insight, we hypothesise that incorporating local information is particularly effective in the context of robot learning, echoing findings in PVR distillation research [39], though this area remains empirically underexplored.
@@
Recognising the importance of local information is only part of the solution. A data-driven mechanism is also required to filter irrelevant details, such as background patches, and prioritise task-relevant information. To this end, we adopt the attentive probing methodology [33] to implement Attentive Feature Aggregation (AFA). Specifically, we append a cross-attention layer to the frozen PVR, modified to include a trainable query token that interacts with the sequence of local tokens produced by the model. These tokens correspond to the per-patch embeddings for ViTs and the channel embeddings for CNNs, both from the final layer.
@@
Consequently, we deploy: $\text{{Attention}}(\boldsymbol{q},F)=\text{{softmax}}\left(\frac{\boldsymbol{q}\cdot(F\cdot W_{K})^{\top}}{\sqrt{d_{k}}}\right)F\cdot W_{V}$. The query token $\boldsymbol{q}$ computes dot products with the feature sequence, with length equal to the number of patches and dimension $d_{k}$, organized as a matrix $F$. These dot products are passed through a softmax function to assign weights to the contributions of each local token to the final embedding. Our AFA module consists of multiple heads, so that specific dimension groups that might be irrelevant to the policy can be filtered out. Gradients are allowed to flow through the cross-attention layer, updating the parameters of $\boldsymbol{q}$ as well as the key and value projection matrices, $W_{K}$ and $W_{V}$.
@@ 4 Experiments
4.1 Implementation Details
Environment. We conduct our experiments in the widely used MetaWorld simulation environment [19], which is built on the MuJoCo [40] physics engine. From this benchmark, we select the ten tasks visualised in Fig. A7 and generate 25 expert demonstrations with a maximum of 175 rollout steps for each using the provided heuristic policies. The primary criterion for task selection is to maintain a balanced representation of easy, medium, and hard tasks, as identified in prior work on PVR-based visuo-motor control [41, 9], as well as from our empirical results.
PVRs. To validate our hypotheses, we deploy seven Residual Networks (ResNets) [42] and seven Vision Transformers (ViT) [43], as summarised in Tab. A1 in the Appendix. Our selection includes the most popular PVRs utilised in robot learning applications that have led to SoTA performance. We also aim to ensure diversity across different training strategies, datasets, and the balance between local and global perception. Despite these variations, we maintain a consistent backbone architecture of ResNet-50 or ViT-B/16, with the exception of DINOv2 [44], which employs a smaller patch size of 14. Also, for DINOv2 we discard overlapping patches, to ensure fairness in the comparison of PVRs. The models tested include powerful representations from vision-specific approaches (e.g., DINO [10]), vision-language models (e.g., CLIP [45]), and robot-learning-focused models (e.g., R3M [27]).
Policy training. For all policy training, we repeat the policy network training five times using different seeds, keeping the PVR frozen, and report the interquartile mean (IQM) success rate. As is common practice in similar work [2, 27, 9], our policy head consists of a small number of MLP layers (four in our case), separated by ReLUs, and outputs the predicted action. We train with mini-batches of 128 samples for 80,000 steps.
Figure 3: Comparison of our Temporal Encoding (TE) against FLARE [25] and using no temporal augmentation on PVR features. Results (sorted by TE) show (a) per-task performance and (b) per-model performance. FLARE and TE bars indicate gains over no temporal information.
@@ 4.2 Temporal Encoding
We evaluate the performance of a policy network trained for each PVR under three conditions: (i) without any temporal component, (ii) using the three most recent past observations and their latent differences (i.e., FLARE [25]), and (iii) with our proposed Temporal Encoding (TE) method. We select the dimensionality of TE to be $64$ and the scale parameter $100$, after tuning these hyperparameters in a subset of tasks and PVRs (see Appendix C.3).
Fig. 3 illustrates (a) the average performance per task for each of the three approaches and (b) the average performance per model. Statistical analysis (paired t-tests, Wilcoxon tests-results in Appendix C.2) confirms that our TE’s gains over FLARE and no augmentation are significant. While VC-1 and iBOT achieve slightly higher average scores with FLARE, all other PVRs benefit significantly from temporal augmentation, even when compared to FLARE-augmented results. Similarly, apart from the “Pick and Place” and “Shelf Place” tasks, TE significantly enhances the average task performance. We attribute TE’s superiority over FLARE to the fact that FLARE’s method of stacking sequential latent embeddings, along with their differences, results in embeddings that are nearly identical due to the high similarity of features extracted from consecutive observations. Consequently, the differences between these embeddings are often near-zero vectors, limiting FLARE’s ability to effectively leverage temporal information. Table A2 provides a detailed summary of the performance of each model-task pair.
An intriguing observation is that PVRs pre-trained with a temporal component in their objective function also benefit considerably from TE. For instance, R3M employs time-contrastive learning [46] to enforce similarity between representations of temporally adjacent frames-experience substantial gains from TE. In addition, VIP achieves an average performance boost of approximately $15$%, making it one of the most positively impacted models. This finding suggests a potential reconsideration of how temporal perception is integrated into features designed for robot learning. It raises the possibility that existing approaches may not fully exploit the temporal structure necessary for optimal performance.
Finally, we also evaluate our TE in conjunction with popular policy modelling techniques that implicitly capture temporal structure. Specifically, we integrate a causal transformer [47, 37] into our policy model and extract features with the strong-performing R3M. This transformer processes a sequence of input tokens (context) and produces a sequence of future actions (horizon). We use rotary positional embeddings [38], which encode the relative position of tokens within the input context, unlike our TE, which encodes the position of each token within the rollout. We apply TE at the output stage of the transformer. To emphasize that TE serves to disentangle temporally adjacent tokens, rather than encode absolute timesteps, we train a multitask policy model across four tasks, using context and horizon length equal to $12$. Even with a single attention layer and only $25$ demonstrations per task, the inclusion of TE led to higher success rates, as shown in Tab. A6.
Figure 4: Robustness evaluation of TE-augmented PVR features, with and without our AFA feature extraction, under lighting and texture perturbations. Subplot (a) reports average performance per task; (b) shows per-model averages.
@@ 4.3 Attentive Feature Aggregation
We use a cross-attention layer with 12 heads for ViT features and 32 heads for ResNet features. This configuration ensures that we process 64-dimensional feature chunks in both cases, maintaining fairness between the two backbone architectures. Consistent with standard attentive probing methodologies [33, 34], we employ a cosine learning rate scheduler with a warm-up phase. TE is applied to augment features with a temporal component and specifically for AFA, the temporal feature is concatenated with the output of the cross-attention module. We validate our full model’s significance via an ablation study, presented in Fig. A4, that evaluates the contributions of TE and AFA in our approach, comparing it against training the policy network solely with PVR-extracted features.
@@
We hypothesise that AFA learns to attend only to scene areas that are important to the task, disregarding features irrelevant to the policy. To properly evaluate this component, we do not limit our evaluation to in-domain scenes but also with environments recreated with scene perturbations, leaving the training dataset distribution unchanged. These perturbations include two modifications, details of which are included in Appendix D.2. First, tabletop texture changes, randomly selected from $30$ distinct textures, some of which feature vibrant patterns that act as strong distractors. These changes affect a significant area of the frame. Second, variations in lighting conditions, including adjustments to the orientation, position, and brightness of the light source. These modifications influence the entire frame, including the robot and the object it manipulates. We summarize the performance of policies trained directly with PVR features and those trained with features aggregated using AFA in Fig. 4, averaged across tasks and models on perturbed scenes. In Fig. A5 we also provide a summary of the performance of each task and model pair on in-domain scenes and on perturbed scenes (per perturbation category).
@@
From these results, several trends emerge. AFA consistently improves the robustness of policies across all tasks, often yielding significant gains, with some tasks showing up to a threefold improvement. The only exception is the “Drawer Open” task, which sees little benefit from AFA. This is likely due to its inherent simplicity, as the task involves manipulating a large object without requiring control of the gripper fingers, which remain open throughout the demonstrations. Consequently, attending to local observations has limited impact. Additionally, most models exhibit improved out-of-distribution performance with AFA, except for CLIP and ViT-B/16. This is reasonable, as these models are trained with objectives that emphasise global frame perception, unlike other models that incorporate supervision at the patch level. Notably, MIM-trained PVRs (i.e., VC-1, DINOv2, MAE and iBOT) benefit the most from AFA, reflecting the alignment between AFA’s design, which is inspired by attentive probing, and the training principles of MIM-based models. These findings highlight AFA’s ability to enhance policy performance, particularly in challenging out-of-distribution scenarios, and underscore its compatibility with models that leverage local feature representations.
@@
The average in-domain performance remains nearly unchanged, with a slight increase from $63.1$% to $66.4$% when using AFA. The minor boost observed with AFA in in-domain scenarios, especially when compared to its substantial improvements in perturbed scenes, suggests that AFA does not learn a new latent space for the PVR, more suited to the task. Instead, it appears to refine the use of the existing latent space by learning to leverage relevant information while discarding elements that are irrelevant to the policy. This distinction underscores AFA’s role as a mechanism for better utilisation of pre-trained features rather than redefining or adapting the underlying feature space.
To further validate that the observed performance boost stems from the use of AFA, and not due to simply pooling local features, we evaluate against two popular alternatives. First, we combine features with Spatial Softmax [15], where we compute the expected 2D coordinates of feature activations across the spatial grid for each channel, effectively summarising the feature map into a set of soft keypoints. Second, we apply TokenLearner [16] configured to extract a single latent token, where a learned spatial attention map is used to selectively aggregate information across all spatial positions into one content-rich representation. We examine both alternatives on all tasks for the top-performing PVR (i.e., VC-1), and we observe (Fig. A10) that AFA leads to much greater out-of-distribution performance, outperforming both by more than 20%.
@@ 5 Discussion
We explored how to effectively utilise PVRs for visuo-motor policy learning by identifying the issues of feature temporal entanglement and lack of robustness in scene visual changes. Furthermore, we proposed two approaches to address these limitations leading to a significant performance increase.
Augmenting PVR-extracted features with TE significantly improved performance, even when combined with SoTA methods that implicitly model temporal structure (i.e., causal transformers with context and action horizons lengths $>1$). We argue that this gain is not due to properties of the heuristic expert demonstrations, which are asynchronous, as Fig. A3 shows. Feature entanglement is also unlikely to be simulation-related, since visualizations in Fig. A1 and Fig. A2 replicate Fig. 1 using real-world demos and reveal similar behaviour. We also evaluated popular video-PVRs, expecting improved temporal modelling. However, as Table A5 indicates, their features were less effective than those from image-PVRs, and still benefited from TE. This counter-intuitive result aligns with the findings of [13], regarding the length of the observation horizon.
Learning to attend to task-related visual cues is imperative for increasing policy robustness under scene perturbations. We argue that this success does not arise from the increase in learning capacity that AFA introduces, since training policy a network using raw PVR features and with a deeper policy network to match AFA’s parameter count did not match AFA’s performance (see Fig. A8). Also, our performance retaliative to TokenLearner and Spatial Softmax further highlights the robustness of AFA under scene perturbations. Additional validation through real-world robot experiments (Section D.6) further supports the generality of our method.
@@ 6 Conclusion
The use of PVRs for visuo-motor policy learning is still in its early stages, and we believe our work highlights key challenges and paves the way for further exploration. In particular, we highlighted characteristics that PVR features currently lack for effective policy learning, such as an inherent sense of task completion and the ability to focus on task-relevant cues. Our insights contribute to the development of PVR models specifically designed for robot learning, ultimately leading to a generalist robotic system powered by large-scale vision foundation models.
Limitations
We conducted extensive experiments, evaluating multiple PVRs and training policies across various seeds and representative tasks. Nonetheless, our study has some limitations.
First, due to the combinatorial scale of the experimental setup—10 tasks, 5 seeds, 14 PVRs, and 4 architectural variants (PVR-only, PVR+TE, PVR+AFA, PVR+TE+AFA), which accumulates to 2,800 training sessions, not including the addition experiments reported in the appendix (e.g., causal transformer, other pooling methods, etc.)—our experiments required significant computational resources (multiple GPU days), limiting our ability to explore a broader range of tasks or other simulation environments.
To mitigate this, we adopted two strategies:
  1. 1.
    We selected challenging MetaWorld tasks that were not trivially solvable. For example, we observed empirically that tasks such as hammer-v2 achieved near-saturated success rates ($\sim$95%) across most PVRs, and thus we omitted it to focus attention on more difficult alternatives.
  2. 2.
    We conducted robustness experiments, similar to the ones in Fig. 4, on a real-world robotic platform to evaluate the transferability of AFA (Section D.6).
Second, while our finding that PVR features exhibit temporal entanglement is significant, its implications in more complex scenarios remain open. In particular, our temporal encoding (TE) approach may struggle to capture meaningful temporal structure in long-horizon tasks involving multiple subgoals or phases. Potentially this problem could be solved by resetting the timestep after each subtask is completed, using a high-level planner to keep track of task progression.
Finally, incorporating local information through AFA introduces a computational trade-off. Since it requires processing sequences of tokens, both training and inference become more resource-intensive, posing challenges for real-time or resource-constrained deployments. Nevertheless, we had no problem running AFA at 10Hz on a NVIDIA GeForce RTX 2080, deploying a trained policy on a KUKA IIWA 14 robot, as discussed in Section D.6.
Acknowledgments
This work was supported by the United Kingdom Research and Innovation (grant EP/S023208/1), EPSRC Centre for Doctoral Training in Robotics and Autonomous Systems (RAS). Funding for DD was provided by ELIAI (Edinburgh Laboratory for Integrated Artificial Intelligence) - EPSRC (EP/W002876/1).
Appendix